Problem 5.16 You invest $270 in a mutual fund today that pays 9.40 percent interest annually. How long will it take to double your money? (If you solve this problem with algebra round intermediate calculations to 6 decimal places, in all cases round your final answer to O decimal place, eg. 545) Number of years
Answer #1

Using financial calculator

PV= -270

I/Y = 9.4%

FV= 270*2 = 540

Find N

N = 7.7153 = 8 years (rounded off)

USing excel, we can find the term as

Sign in - Book1 - Excel 《Product Activation Failed) Share Tell me what you want to do Home InsertDrawPage Layout FormulasData
Answer #2

SOLUTION :


FV / PV = (1 + r)^n

=> 2 = (1 + 0.094)^n = 1.094^n 

Taking log :

log(2) = n log(1.094)

=> n = log(2)/log(1.094)

=> n = 7.714 years (ANSWER) 

answered by: Tulsiram Garg

> Rounded off to 0 decimals = 8.0 % (ANSWER)

Tulsiram Garg Tue, Mar 29, 2022 1:14 PM